# Domain-Aware Tensor Network Structure Search

### v1

@@ 4.1 Accelerating SOTA methods with tnLLM
In the first part of Section 4, we have demonstrated that tnLLM achieves comparable objective function values compared to SOTA sampling-based algorithms, while generating domain-aware explanations for the identified TN structures and significantly reducing the required number of function evaluations. However, due to the black-box nature of LLMs, no theoretical analysis can be provided for the evaluation efficiency of tnLLM, in contrast to TNLS [11] and TnALE [10].
At the same time, while sampling-based methods follow a “local-search” scheme within a neighborhood, there is no guarantee on the minimum number of evaluations required to find a “good” neighborhood. Moreover, poor initialization significantly increases the number of evaluations needed. To address this, we construct a hybrid algorithm that first runs tnLLM for 10 evaluations to leverage its domain knowledge and reasoning capabilities in order to identify a strong initialization point in the “global-search” stage. Sampling-based methods then perform “local-search” in the identified neighborhood, combining the speed-up benefits of tnLLM with the theoretical guarantees of sampling-based approaches.
@@
Figure 7 demonstrates the performance difference of TNLS and TnALE with and without a tnLLM-initialized structure. The “global-search” capability of tnLLM significantly sped up the iterative minimization process. During “local-search”, both TNLS and TnALE further improved upon the structures found in “global-search” by tnLLM. Overall, the tnLLM-initialized algorithms achieved nearly identical objective function values with up to 23×23$\times$23 × fewer evaluations compared to vanilla TNLS, and up to 13×13$\times$13 × fewer evaluations compared to vanilla TnALE. The full numerical performance comparison is provided in Appendix F.
@@ 4.2 Ablation study
To assess the effectiveness of injecting domain information into the TN-SS problem, we removed the carefully designed structured prompts of the tnLLM framework, equivalent to using the LLM to solve the TN-SS problem without any priors on the domain information. Without domain information, the LLM generates a poorly initialized TN structure that is 80.3% worse in terms of objective function value, due to its lack of information about the different modes, and requires 10.4×10.4$\times$10.4 × more evaluations to converge. Moreover, as the model defaults to random assumptions about the modes, as illustrated in Figure 6, it was found to produce uniformly connected TN structures throughout the minimization process. This restricts the search space to a much smaller set of candidate solutions that are not domain-meaningful from a practical perspective.
@@
To examine how the choice of LLM affects the performance of tnLLM, we compared our baseline model, GPT-4o (gpt-4o-2024-08-06), against GPT-4.5 (gpt-4.5-preview-2025-02-27), GPT4o-mini (gpt-4o-mini-2024-07-18) and GPT-3.5 (gpt-3.5-turbo-1106). Observe from Table 2 that due to the carefully designed prompts and the structure of the overall framework, tnLLM is robust to the choice of the LLM used. While there are small variations in the achieved objective function values, the overall performance remained consistent across all three datasets in both training and test sets. Furthermore, the number of evaluations required by all LLMs falls within the mean ± standard deviation range reported in Table 1. It is important to note, however, that weaker models, and in particular GPT-3.5, are more prone to “hallucinations”. For example, we observed instances where the model misinterpreted the optimization objective, mistakenly assuming that higher objective values were always preferable.
@@
Dataset GPT-4o (Baseline) GPT-4.5 GPT-4o-mini GPT-3.5
Train Test Train Test Train Test Train Test
Images -0.63 -0.48 -0.62 -0.54 -0.61 -0.48 -0.64 -0.52
Videos -1.63 -1.70 -1.62 -1.65 -1.62 -1.69 -1.63 -1.70
Time-series -0.42 -0.41 -0.40 -0.41 -0.41 -0.40 -0.42 -0.40
@@ Appendix A Limitations and future work
Despite the good performance, the proposed domain-aware TN-SS framework currently lacks theoretical guarantees for its evaluation efficiency due to the black-box nature of LLMs. This could be mitigated by developing domain-aware TN-SS algorithms based on continuous optimization. Also, as mentioned in the ablation study, weaker LLM models, such as GPT-3.5, were found to be more prone to misinterpret the optimization objectives. Therefore, improving the reasoning consistency of smaller LLM models promises to further improve the efficiency of the framework.
@@ Appendix C Data
We constructed a custom financial time series dataset with 142 temporally ordered fifth-order tensors, denoted as $\{\mathcal{X}_{n}\}_{n=1}^{142}\in\mathbb{R}^{3\times 6\times 3\times 4\times 5}$, each representing a rolling window produced via multi-way delay embedding through the temporal direction [38]. This leads to the value selection process of $10$ ranks. To the best of our knowledge, this is the largest tensor dataset in terms of number of samples ever considered in the TN-SS problem. The first $80\%$ of these tensors were used as the training data, while the remaining $20\%$ with non-overlapping entries with the training data were used for testing. The modes of each time series tensor correspond to:
@@
Video dataset is produced from performing the multi-way delay embedding [6] through the temporal mode in a video to produce 5 samples, $\{\mathcal{X}_{n}\}_{n=1}^{5}\in\mathbb{R}^{144\times 176\times 3\times 10}$. Figure 12 shows an example sample of the videos dataset at different frames. The modes of each video tensor correspond to:
@@ Appendix D Implementation details
All baseline models are initialized as a “fully-disconnected” graph. The TN search template is set to a complete graph. For TNLS, we set the number of samples (evaluations), #$\#Sample$, in each sampling step to $4$ in the images dataset, $60$ in the videos dataset, and $10$ in the time series dataset. For TnALE, we set $D=1$, $r=1$. For tnGPS, we selected the best-performing algorithm discovered (“Ho-2”) and adopted the following hyperparameters as suggested in their paper: code upper bound of 10, mutation rate of 0.1, crossover rate of 0.6, selection pressure of 1.5, elitism enabled, diversity factor of 0.05, variance decay of 0.98, minimum variance of 0.1, tournament size factor of 0.2, elite diversity boost of 2.0, random individual chance of 0.05, and a maximum mutation of 3. All models were run for $500$ evaluations in both the images and videos dataset, and $250$ evaluations in the time series dataset if not converged. An internal server with NVIDIA RTX A6000 GPU, an AMD Ryzen Threadripper PRO 5955WX with 16 cores, and 256GB of RAM was used.
Appendix E Summarized generated explanations of tnLLM for the initialized TN structure
The full set of summarized LLM generated explanations over three independent runs for the images, videos, and time-series datasets are provided in Tables 3, 4, and 5, respectively. The suggested ranks are verified to be consistent with the domain-informed reasoning provided by tnLLM. Also, the domain-relevant explanations are robust across multiple runs.
@@ Appendix E Summarized generated explanations of tnLLM for the initialized TN structure
Appendix F Performance comparison for vanilla and tnLLM-initialized TNLS and TnALE
@@
Table 6: Performance comparison across different datasets for vanilla and tnLLM-initialized TNLS and TnALE. The values on the left give the lowest training objective function values and their corresponding testing objective function values. The values in [square brackets] give the number of evaluations to first achieve the best training objective function value. For both metrics, a lower value is better. The least number of evaluations are highlighted in bold.
@@
Time-series Train -0.45 [218] -0.47 [14] -0.47 [177] -0.47 [20]
@@
Appendix G Broader societal impact
This work uses LLMs to directly infer tensor network (TN) structures without human intervention by incorporating domain information about the data and utilizing the inherent reasoning capabilities of LLMs. The proposed framework significantly enhances efficiency and reduces computational costs compared to traditional methods, including sampling-based approaches which to date have demonstrated superior performance. By lowering computational demands, our approach also reduces $\text{CO}_{2}$ emissions and the environmental impact, a growing concern with big-data applications that often involve tensors. Furthermore, our approach provides domain-relevant explanations for the identified TN structures, enabling domain experts to verify the identified TN structures. This is particularly crucial in sensitive domains such as finance and healthcare.

### v2

@@ 4.1 Accelerating SOTA methods with tnLLM
In the first part of Section 4, we have demonstrated that tnLLM achieves comparable objective function values compared to SOTA sampling-based algorithms, while generating domain-aware explanations for the identified TN structures and significantly reducing the required number of function evaluations. However, due to the black-box nature of LLMs, no theoretical analysis can be provided for the evaluation efficiency of tnLLM, in contrast to TNLS and TnALE.
At the same time, while sampling-based methods follow a “local-search” scheme within a neighborhood, there is no guarantee on the minimum number of evaluations required to find a “good” neighborhood. Moreover, poor initialization significantly increases the number of evaluations needed. To address this, we construct a hybrid algorithm that first runs tnLLM for 10 evaluations to leverage its domain knowledge and reasoning capabilities in order to identify a strong initialization point in the “global-search” stage. Sampling-based methods then perform “local-search” in the identified neighborhood, combining the speed-up benefits of tnLLM with their theoretical guarantees.
@@
Figure 7 demonstrates the performance difference of TNLS and TnALE with and without a tnLLM-initialized structure. The “global-search” capability of tnLLM significantly sped up the iterative minimization process. During “local-search”, both TNLS and TnALE further improved upon the structures found in “global-search” by tnLLM. Overall, the tnLLM-initialized algorithms achieved nearly identical objective function values with up to $23\times$ fewer evaluations compared to vanilla TNLS, and up to $13\times$ fewer evaluations compared to vanilla TnALE. The full numerical performance comparison is provided in Appendix G.
@@ 4.2 Ablation study
To assess the effectiveness of injecting domain information into the TN-SS problem, we removed the carefully designed structured prompts of the tnLLM framework, equivalent to using the LLM to solve the TN-SS problem without any priors on the domain information. Without domain information, the LLM generates a poorly initialized TN structure that is 80.3% worse in terms of objective function value, due to its lack of information about the different modes, and requires $10.4\times$ more evaluations to converge. Moreover, as the model defaults to random assumptions about the modes, as illustrated in Figure 6, it was found to produce uniformly connected TN structures throughout the minimization process. This restricts the search space to a much smaller set of solutions that are not domain-meaningful from a practical perspective.
@@
To examine how the choice of LLM affects the performance of tnLLM, we compared our baseline model, GPT-4o (gpt-4o-2024-08-06), against GPT-4.5 (gpt-4.5-preview-2025-02-27), GPT4o-mini (gpt-4o-mini-2024-07-18), GPT-3.5 (gpt-3.5-turbo-1106) and the open-source DeepSeek V3 (DeepSeek-V3-0324) model. Observe from Table 2 that due to the carefully designed prompts and the structure of the overall framework, tnLLM is robust to the choice of the LLM used. While there are small variations in the achieved objective function values, the overall performance remained consistent across all three datasets in both training and testing sets. Furthermore, the number of evaluations required by all LLMs falls within the mean ± standard deviation range reported in Table 1. It is important to note that, despite the consistent performance, weaker models and in particular GPT-3.5 occasionally misinterpreted the optimization objective, mistakenly assuming that higher values were preferable.
@@
Dataset GPT-4o (Baseline) GPT-4.5 GPT-4o-mini GPT-3.5 DeepSeek-V3
Train Test Train Test Train Test Train Test Train Test
Images -0.63 -0.48 -0.62 -0.54 -0.61 -0.48 -0.64 -0.52 -0.62 -0.47
Videos -1.63 -1.70 -1.62 -1.65 -1.62 -1.69 -1.63 -1.70 -1.63 -1.70
Time-series -0.42 -0.41 -0.40 -0.41 -0.41 -0.40 -0.42 -0.40 -0.40 -0.39
@@ Appendix A Limitations and future work
Despite the good performance, the proposed domain-aware TN-SS framework currently lacks theoretical guarantees for its evaluation efficiency due to the black-box nature of LLMs. This could be mitigated by developing domain-aware TN-SS algorithms based on continuous optimization. Also, as mentioned in the ablation study, weaker LLM models, such as GPT-3.5, were found to be more prone to misinterpret the optimization objectives. Therefore, improving the reasoning consistency of smaller LLM models promises to further improve the efficiency of the framework. Furthermore, prompt learning is an active area of research in LLMs, which can be used to potentially enhance the tnLLM framework in future works.
@@ Appendix C Data
We constructed a custom financial time series dataset with 142 temporally ordered fifth-order tensors, denoted as $\{\mathcal{X}_{n}\}_{n=1}^{142}\in\mathbb{R}^{3\times 6\times 3\times 4\times 5}$, each representing a rolling window produced via multi-way delay embedding through the temporal direction (Yokota et al., 2018). This leads to the value selection process of $10$ ranks. To the best of our knowledge, this is the largest tensor dataset in terms of number of samples ever considered in the TN-SS problem. The first $80\%$ of these tensors were used as the training data, while the remaining $20\%$ with non-overlapping entries with the training data were used for testing. The modes of each time series tensor correspond to:
@@
Video dataset is produced from performing the multi-way delay embedding (Yamamoto et al., 2022) through the temporal mode in a video to produce 5 samples, $\{\mathcal{X}_{n}\}_{n=1}^{5}\in\mathbb{R}^{144\times 176\times 3\times 10}$. Figure 12 shows an example sample of the videos dataset at different frames. The modes of each video tensor correspond to:
@@ Appendix D Implementation details
All baseline models are initialized as a “fully-disconnected” graph. The TN search template is set to a complete graph. For TNLS, we set the number of samples (evaluations), #Sample\#Sample, in each sampling step to $4$ in the images dataset, $60$ in the videos dataset, and $10$ in the time series dataset. For TnALE, we set $D=1$, $r=1$. For tnGPS, we selected the best-performing algorithm discovered (“Ho-2”) and adopted the following hyperparameters as suggested in their paper: code upper bound of 10, mutation rate of 0.1, crossover rate of 0.6, selection pressure of 1.5, elitism enabled, diversity factor of 0.05, variance decay of 0.98, minimum variance of 0.1, tournament size factor of 0.2, elite diversity boost of 2.0, random individual chance of 0.05, and a maximum mutation of 3. All models were run for $500$ evaluations in both the images and videos dataset, and $250$ evaluations in the time series dataset if not converged. An internal server with NVIDIA RTX A6000 GPU, an AMD Ryzen Threadripper PRO 5955WX with 16 cores, and 256GB of RAM was used.
Appendix E Runtime Comparisons
Table 3 reports the total average runtime in seconds to first achieve the best training objective function value for all models across the three datasets of Table 1. For tnLLM, this also includes the LLM inference time. The proposed tnLLM achieves comparable performance to SOTA methods while reducing runtime by up to 98.3% compared to TNLS, 97.7% compared to TnALE, and 98.1% compared to tnGPS.
Table 3: Runtime comparisons across three datasets. The values are the total average runtime in seconds to first achieve the best training objective function value. A lower value is better. The best values are denoted in bold. The second best values are underlined.
Data Type TNLS TnALE tnGPS tnLLM (Ours)
Images 2,122s 1,659s 8,025s 151s
Videos 17,690s 15,298s 20,461s 1,984s
Time-series 110,945s 80,700s 17,287s 1,866s
Appendix F Summarized generated explanations of tnLLM for the initialized TN structure
The full set of summarized LLM generated explanations over three independent runs for the images, videos, and time-series datasets are provided in Tables 4, 5, and 6, respectively. The suggested ranks are verified to be consistent with the domain-informed reasoning provided by tnLLM. Also, the domain-relevant explanations are robust across multiple runs.
@@ Appendix F Summarized generated explanations of tnLLM for the initialized TN structure
Appendix G Performance comparison for vanilla and tnLLM-initialized TNLS and TnALE
@@
Table 7: Performance comparison across different datasets for vanilla and tnLLM-initialized TNLS and TnALE. The values on the left give the lowest training objective function values and their corresponding testing objective function values. The values in [square brackets] give the number of evaluations required to first achieve the best training objective function value. For both metrics, a lower value is better. The least number of evaluations are highlighted in bold.
@@
Time‐series Train -0.45 [218] -0.47 [14] -0.47 [177] -0.47 [20]